# Scaling Limits of Line Models in Degenerate Environment

### v1

@@ 4.2.4 A time change
We also denote by $Y^{T}$ the process with generator as in (4.10) but with the random variables $H$ replaced by $H^{(T)}$ as defined in (4.2). The advantage of considering $(Y(t))_{t\geq 0}$, is that, for all $T>0$, the vertical component $(Y^{T}_{2}(t))_{t\geq 0}$ is now a simple random walk independent of the environment. In particular, convergence holds for its diffusively rescaled local times thanks to Lemma 5.6 (see Appendix).
@@
$\langle Y_{1}^{T},Y_{1}^{T}\rangle_{t}=\int_{0}^{t}\frac{H^{(T)}(Y_{2}^{T}(s))}{V(Y_{1}^{T}(s))}ds,\qquad t\geq 0.$
@@
$\mu^{T}(dx)=\frac{1}{\sqrt{T}^{1/\alpha_{1}}}\sum_{x\in 1/\sqrt{T}\mathbb{Z}}H(\sqrt{T}x)\delta_{x}=\sum_{x\in T^{-1/2}\mathbb{Z}}\left(\mathcal{H}\left(x+\frac{1}{\sqrt{T}}\right)-\mathcal{H}(x)\right)\,\delta_{x}.$
@@
$\langle Y^{(T)}_{1},Y^{(T)}_{1}\rangle_{t}=\frac{1}{T^{\delta}}\int_{0}^{Tt}\frac{H^{(T)}(Y^{T}_{2}(s))}{V(Y_{1}^{T}(s))}\,ds=D^{T}(t)$
which converges in law, when $T\to\infty$, to $(\Delta(t))_{t\geq 0}$. An important fact in the sequel will be the fact that the martingale property is preserved when we take a limit in law. This will be guaranteed by the following lemma:
@@
$\int_{\mathbb{R}}\int_{\mathbb{R}}E^{\omega}_{0}[\ell^{Y,T}_{1}(x)^{4}]^{1/4}\,E^{\omega}_{0}[\ell^{Y,T}_{1}(y)^{4}]^{1/4}\,\mathbb{P}(\max_{[0,T]}|Y_{2}|\geq|x|)\,\mathbb{P}(\max_{[0,T]}|Y_{2}|\geq|y|)\,d\mu^{T}(x)\,d\mu^{T}(y).$
@@
$\begin{split}\mathbb{P}(\max_{[0,T]}|Y_{2}|\geq\sqrt{T}|x|)&=\sum_{n=0}^{\infty}\mathbb{P}\left(N_{T}=n,\max_{0\leq j\leq n}|S_{j}|\geq|x|\sqrt{T/n}\right)\\ &\leq cT^{-k}\sum_{n=0}^{\infty}\mathbb{P}(N_{T}=n)\,n^{k}|x|^{-2k}=cT^{-k}\,\mathbb{E}[N_{T}^{k}]\,|x|^{-2k}.\end{split}$
@@
indeed this follows easily upon writing
@@ 4.2.5 Martingale convergence using a Girsanov argument
When $T\to\infty$, the process $(Y^{(T)}(t))_{t\geq 0}$ converges in law, in $D([0,\infty),\mathbb{R})^{2}$, to the process $(\mathcal{Y}_{1}(t),\mathcal{Y}_{2}(t))_{t\geq 0}$, where
@@
$\langle Y^{(T)}_{1},Y^{(T)}_{1}\rangle_{t}=2D^{T}(t),\quad\langle Y^{(T)}_{2},Y^{(T)}_{2}\rangle_{t}=2t,\quad\langle Y^{(T)}_{1},Y^{(T)}_{2}\rangle_{t}=0\qquad t\geq 0,$
where we recall that $(D^{T}(t))_{t\geq 0}$ converges in law, as $T\to\infty$, to the Kesten-Spitzer process $(\Delta(t))_{t\geq 0}$. Note further that $Y^{(T)}_{2}=\frac{1}{\sqrt{T}}Y_{2}(Tt)$ where $Y_{2}$ is a simple random walk, while we can write $Y^{(T)}_{1}(t)=S^{T^{\delta}}_{1}\circ D^{T}$, where
@@
$Y^{(T)}_{1}(t),\quad Y^{(T)}_{2}(t),\quad Y^{(T)}_{1}(t)^{2}-2D^{T}(t),\quad Y^{(T)}_{1}(t)Y^{(T)}_{2}(t),\quad\hbox{and}\quad Y^{(T)}_{2}(t)-2t$
@@ 4.2.6 Conclusion for the original model
From the relation (4.11), we have, for all $t\geq 0$,
$\int_{0}^{t}H(X_{2}(s))\,ds=\int_{0}^{A(t)}\frac{H(Y_{2}(r))}{V(Y_{1}(r))}\,dr.$
Here we performed the change of variable $r=A(s)$ to obtain the right-hand side. It therefrom follows that
$\frac{1}{T^{\delta}}\int_{0}^{Tt}H(X_{2}(s))\,ds=D^{V,T}\circ A^{(T)}(t),$
where $D^{V,T}$ was defined in (4.12), and where, for all $t\geq 0$, $A^{(T)}(t):=\frac{1}{T}A(Tt)$. Now we have proven above that $D^{V,T}\underset{T\to\infty}{\Longrightarrow}\Delta$ in $C$. In addition we have the convergence of processes
$A^{(T)}(t):=\frac{1}{T}\int_{0}^{Tt}V(X_{1}(s))\,ds\underset{T\to\infty}{\longrightarrow}t$
@@
$(X^{(T)}_{1},X^{(T)}_{2},D^{T})\underset{T\to\infty}{\Longrightarrow}(B_{1}\circ\Delta,B_{2},\Delta).$ (4.17)
where $B_{1},B_{2},\Delta$ are as in the statement of Theorem 4.3. This enhanced convergence will be useful in the next section.
@@ 4.2.7 Scaling limit of the Constant Speed Random Walk
$J^{(M)}(t)=D^{M}(t)+\frac{1}{M^{\delta}}\int_{0}^{Mt}V(X_{1}^{T}(s))\,ds,\qquad t\geq 0$
where $D^{M}(t)$ was defined in (4.12). By (4.4), $D^{M}\underset{T\to\infty}{\Longrightarrow}\Delta$ in $D$; moreover, since $\delta>1$, there holds the convergence of processes, in $C$,
$\frac{1}{M^{\delta}}\int_{0}^{Mt}V(X_{1}(s))\,ds\underset{M\to\infty}{\longrightarrow}0\qquad\hbox{a.s.}$
@@
$(X^{(T^{1/\delta})}_{1},X^{(T^{1/\delta})}_{2},I^{(T^{1/\delta})})\underset{T\to\infty}{\Longrightarrow}(B_{1}\circ\Delta,B_{2},\Delta).$
@@
$(X^{(T^{1/\delta})}_{1},X^{(T^{1/\delta})}_{2},{I^{(T^{1/\delta})}}^{-1})\underset{T\to\infty}{\Longrightarrow}(B_{1}\circ\Delta,B_{2},\Delta^{-1}),$
@@ 5.2 A useful estimate for maxima of heavy-tailed random variables
Let $\alpha\in(0,1)$ and $Z_{k}$, $k\in\mathbb{Z}$, be a collection of i.i.d. random variables defined on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$ with tail behavior $\mathbb{P}(Z_{k}>t)=t^{-\alpha+o(1)}$ as $t\to\infty$, where $\alpha>0$. Let $\alpha^{-}\in(0,\alpha)$. Then, for $\mathbb{P}$-a.e. $\omega$, there exists $C(\omega)>0$, such that, for all $k\geq 1$,
@@ 5.3 Estimates on solutions of a system of coupled non-linear differential inequalities
Let $A_{i}:=\frac{1+\varepsilon_{i}^{+}}{1-\varepsilon_{1}^{+}\varepsilon_{2}^{+}}$, $i=1,2$. Let $C,h_{0},v_{0}>0$, and let $h,v:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be two continuous functions.
@@
    $\forall t\in[0,T],\quad\begin{cases}h(t)\leq C(h_{0}+A_{1}\int_{0}^{t}v(s)^{\varepsilon_{1}^{+}}\,ds)\\ v(t)\leq C(v_{0}+A_{2}\int_{0}^{t}h(s)^{\varepsilon_{2}^{+}}\,ds),\end{cases}$ (5.2)
@@
    $\forall t\in[0,T],\quad\begin{cases}h(t)\leq C(h_{0}+A_{1}\int_{0}^{t}v(s)^{\varepsilon_{1}^{+}\varepsilon_{2}^{+}}\,ds)\\ v(t)\leq C(v_{0}+\frac{A_{2}}{\varepsilon_{2}^{+}}\int_{0}^{t}h(s)^{\varepsilon_{2}^{+}}v(s)^{1-\varepsilon_{2}^{+}}\,ds),\end{cases}$ (5.3)
@@
$x(t)=[(C^{+}h_{0})^{1/A_{1}}+C^{+}t]^{A_{1}},\quad y(t)=[(C^{+}v_{0})^{1/A_{2}}+C^{+}t]^{A_{2}},\qquad t\geq 0.$
@@
$\forall t\in[0,T],\quad\begin{cases}x(t)=C^{+}(h_{0}+A_{1}\int_{0}^{t}(y(s))^{\varepsilon_{1}^{+}}\,ds)\\ y(t)=C^{+}(v_{0}+A_{2}\int_{0}^{t}(x(s))^{\varepsilon_{2}^{+}}\,ds).\end{cases}$ (5.4)
@@
$h(t)<C(h_{0}+A_{1}\int_{0}^{t}(y(s))^{\varepsilon_{1}^{+}}\,ds)<C^{+}(h_{0}+A_{1}\int_{0}^{t}(y(s))^{\varepsilon_{1}^{+}}\,ds)=x(t)$
@@
$\forall t\in[0,T],\quad h(t)<[(C^{+}h_{0})^{1/A_{1}}+C^{+}t]^{A_{1}}\quad\hbox{and}\quad v(s)<[(C^{+}v_{0})^{1/A_{2}}+C^{+}t]^{A_{2}},$
@@
$x(t)=[(C^{+}h_{0})^{1/A_{1}}+C^{+}t)]^{A_{1}},\quad y(t)=[(C^{+}v_{0})^{1/B_{2}}+C^{+}t]^{B_{2}},\qquad t\geq 0,$
@@
$\forall t\in[0,T],\quad\begin{cases}x(t)=C^{+}(h_{0}+A_{1}\int_{0}^{t}y(s)^{\varepsilon_{1}^{+}\varepsilon_{2}^{+}}\,ds)\\ y(t)=C^{+}(v_{0}+B_{2}\int_{0}^{t}x(s)^{\varepsilon_{2}^{+}}y(s)^{1-\varepsilon_{2}^{+}}\,ds).\end{cases}$ (5.5)
@@ 5.4 Convergence of local times of a continuous-time simple random walk
for all $t,s\in[0,M]$ and $x,y\in\mathbb{R}$. Since $\tilde{\ell}^{T}_{t}(\cdot)$ is the piecewise linear interpolation of $\ell^{T}_{t}(\cdot)$, by a standard argument, equation (5.7) will follow upon proving
@@
Note that $\left(\frac{1}{\sqrt{T}}\xi(Tt)\right)_{t\geq 0}$ and $\left(\frac{1}{\sqrt{T}}M_{f_{\sqrt{T}x}}(Tt)\right)_{t\geq 0}$ are both martingales with q.v. bounded by $\frac{1}{T}Tt=t$, so, by the BDG inequality, both expectations above are bounded by $t^{p}$, up to a constant depending only on $p$. Therefore we obtain the following bound which is uniform in $x\in\frac{1}{\sqrt{T}}\mathbb{Z}$:
@@
where $(M_{f_{k,j}}(t))_{t\geq 0}$ is a martingale with q.v. $\langle M_{f_{k,j}},M_{f_{k,j}}\rangle_{t}\leq 2\int_{0}^{t}\mathbf{1}_{\{k\leq\xi(s)\leq j\}}\,ds$. We deduce that, for all $x,y\in\frac{1}{\sqrt{T}}\mathbb{Z}$ with $x<y$,
@@
where the martingale $M^{T}(t):=\frac{1}{\sqrt{T}}M_{f_{\sqrt{T}x,\sqrt{T}y}}(tT)$, $t\geq 0$, has q.v.
@@
Thus, using the triangle inequality and the inequality $(a+b)^{p}\leq 2^{p}(a^{p}+b^{p})$, we get

### v2

@@ 4.2.4 A time change
We also denote by $Y^{T}$ the process with generator as in (4.10) but with the random variables $H$ replaced by $H^{(T)}$ as defined in (4.2). The advantage of performing this time change is that, for all $T>0$, the vertical component $(Y^{T}_{2}(t))_{t\geq 0}$ is now a simple random walk independent of the environment. In particular, convergence holds for its diffusively rescaled local times thanks to Lemma 5.6 (see Appendix).
@@
$\langle Y_{1}^{T},Y_{1}^{T}\rangle_{t}=2\int_{0}^{t}\frac{H^{(T)}(Y_{2}^{T}(s))}{V(Y_{1}^{T}(s))}ds,\qquad t\geq 0.$
@@
$\mu^{T}(dx)=\frac{1}{T^{\frac{1}{2\alpha_{1}}}}\sum_{x\in T^{-1/2}\mathbb{Z}}H(\sqrt{T}x)\delta_{x}=\sum_{x\in T^{-1/2}\mathbb{Z}}\left(\mathcal{H}\left(x+\frac{1}{\sqrt{T}}\right)-\mathcal{H}(x)\right)\,\delta_{x}.$
@@
$\langle Y^{(T)}_{1},Y^{(T)}_{1}\rangle_{t}=\frac{2}{T^{\delta}}\int_{0}^{Tt}\frac{H^{(T)}(Y^{T}_{2}(s))}{V(Y_{1}^{T}(s))}\,ds=2\,D^{V,T}(t)$
which converges in law, when $T\to\infty$, to $(2\,\Delta(t))_{t\geq 0}$. An important fact in the sequel will be the fact that the martingale property is preserved when we take a limit in law. This will be guaranteed by the following lemma:
@@
$\int_{\mathbb{R}^{2}}E^{\omega}_{0}[\ell^{Y,T}_{1}(x)^{4}]^{\frac{1}{4}}\,E^{\omega}_{0}[\ell^{Y,T}_{1}(y)^{4}]^{\frac{1}{4}}\,\mathbb{P}(\max_{[0,T]}|Y_{2}|\geq\sqrt{T}|x|)\,\mathbb{P}(\max_{[0,T]}|Y_{2}|\geq\sqrt{T}|y|)\,d\mu^{T}(x)\,d\mu^{T}(y).$
@@
$\begin{split}\mathbb{P}(\max_{[0,T]}|Y_{2}|\geq\sqrt{T}|x|)&=\sum_{n=0}^{\infty}\mathbb{P}\left(N_{T}=n,\max_{0\leq j\leq n}|S_{j}|\geq|x|\sqrt{\frac{T}{n}}\sqrt{n}\right)\\ &\leq cT^{-k}\sum_{n=0}^{\infty}\mathbb{P}(N_{T}=n)\,n^{k}|x|^{-2k}=cT^{-k}\,\mathbb{E}[N_{T}^{k}]\,|x|^{-2k}.\end{split}$
@@
as can be verified upon writing
@@ 4.2.5 Martingale convergence using a Girsanov argument
For $\mathbb{P}$ a.e. environment, when $T\to\infty$, the process $(Y^{(T)}(t))_{t\geq 0}$ converges in law, in $D([0,\infty),\mathbb{R})^{2}$, to the process $(\mathcal{Y}_{1}(t),\mathcal{Y}_{2}(t))_{t\geq 0}$, where
@@
$\langle Y^{(T)}_{1},Y^{(T)}_{1}\rangle_{t}=2D^{V,T}(t),\quad\langle Y^{(T)}_{2},Y^{(T)}_{2}\rangle_{t}=2t,\quad\langle Y^{(T)}_{1},Y^{(T)}_{2}\rangle_{t}=0\qquad t\geq 0,$
where we recall that $(D^{V,T}(t))_{t\geq 0}$ converges in law, as $T\to\infty$, to the Kesten-Spitzer process $(\Delta(t))_{t\geq 0}$. Note further that $Y^{(T)}_{2}=\frac{1}{\sqrt{T}}Y_{2}(Tt)$ where $Y_{2}$ is a simple random walk, while we can write $Y^{(T)}_{1}=S^{T^{\delta}}_{1}\circ D^{T}$, where
@@
$Y^{(T)}_{1}(t),\quad Y^{(T)}_{2}(t),\quad Y^{(T)}_{1}(t)^{2}-2D^{V,T}(t),\quad Y^{(T)}_{1}(t)Y^{(T)}_{2}(t),\quad\hbox{and}\quad Y^{(T)}_{2}(t)-2t$
@@ 4.2.6 Conclusion for the original model
From the relation (4.11), we have, for all $T>0$,
$\int_{0}^{t}H(X^{T}_{2}(s))\,ds=\int_{0}^{A(t)}\frac{H(Y^{T}_{2}(r))}{V(Y^{T}_{1}(r))}\,dr,\qquad\hbox{where}\qquad A(t)=\int_{0}^{t}V(X^{T}_{1}(s))ds,\quad t\geq 0.$
In the first equality we performed the change of variable $r=A(s)$ to obtain the right-hand side. It therefrom follows that, setting
$A^{H,T}_{1}(t):=\frac{1}{T^{\delta}}\int_{0}^{Tt}H(X_{2}^{T}(s))\,ds,$ (4.17)
then $A^{H,T}_{1}=D^{V,T}\circ A^{(T)}$, where $D^{V,T}$ was defined in (4.12), and where, for all $t\geq 0$, $A^{(T)}(t):=\frac{1}{T}A(Tt)$. Now we have proven above that $D^{V,T}\underset{T\to\infty}{\Longrightarrow}\Delta$ in $C$. In addition we have the convergence of processes
$A^{(T)}(t):=\frac{1}{T}\int_{0}^{Tt}V(X^{T}_{1}(s))\,ds\underset{T\to\infty}{\longrightarrow}t$
@@
$(X^{(T)}_{1},X^{(T)}_{2},A^{H,T}_{1})\underset{T\to\infty}{\Longrightarrow}(B_{1}\circ\Delta,B_{2},\Delta).$ (4.18)
where $A^{H,T}_{1}$ is as in (4.17), and where $B_{1},B_{2},\Delta$ are as in the statement of Theorem 4.3. This enhanced convergence will be useful in the next section.
@@ 4.2.7 Scaling limit of the Constant Speed Random Walk
$J^{(M)}(t)=A^{H,M}_{1}(t)+\frac{1}{M^{\delta}}\int_{0}^{Mt}V(X_{1}^{T}(s))\,ds,\qquad t\geq 0$
where $A^{H,M}_{1}(t)$ was defined in (4.17). By (4.4), $A^{H,M}_{1}(t)\underset{T\to\infty}{\Longrightarrow}\Delta$ in $D$; moreover, since $\delta>1$, there holds the convergence of processes, in $C$,
$\frac{1}{M^{\delta}}\int_{0}^{Mt}V(X^{T}_{1}(s))\,ds\underset{M\to\infty}{\longrightarrow}0\qquad\hbox{a.s.}$
@@
$(X^{(T^{1/\delta})}_{1},X^{(T^{1/\delta})}_{2},J^{(T^{1/\delta})})\underset{T\to\infty}{\Longrightarrow}(B_{1}\circ\Delta,B_{2},\Delta).$
@@
$(X^{(T^{1/\delta})}_{1},X^{(T^{1/\delta})}_{2},{J^{(T^{1/\delta})}}^{-1})\underset{T\to\infty}{\Longrightarrow}(B_{1}\circ\Delta,B_{2},\Delta^{-1}),$
@@ 5.2 A useful estimate for maxima of heavy-tailed random variables
Let $\alpha\in(0,1)$ and $Z_{k}$, $k\in\mathbb{Z}$, be a collection of i.i.d. random variables defined on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$ with tail behavior $\mathbb{P}(Z_{k}>t)=t^{-\alpha+o(1)}$ as $t\to\infty$. Let $\alpha^{-}\in(0,\alpha)$. Then, for $\mathbb{P}$-a.e. $\omega$, there exists $C(\omega)>0$, such that, for all $k\geq 1$,
@@ 5.3 Estimates on solutions of a system of coupled non-linear differential inequalities
Let $A_{i}:=\frac{1+\varepsilon_{i}^{+}}{1-\varepsilon_{1}^{+}\varepsilon_{2}^{+}}$, $i=1,2$. Let $C,\kappa>0$, and let $h,v:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be two continuous functions.
@@
    $\forall t\in[0,T],\quad\begin{cases}h(t)\leq C(\kappa+A_{1}\int_{0}^{t}v(s)^{\varepsilon_{1}^{+}}\,ds)\\ v(t)\leq C(\kappa+A_{2}\int_{0}^{t}h(s)^{\varepsilon_{2}^{+}}\,ds),\end{cases}$ (5.2)
@@
    $\forall t\in[0,T],\quad\begin{cases}h(t)\leq C(\kappa+A_{1}\int_{0}^{t}v(s)^{\varepsilon_{1}^{+}\varepsilon_{2}^{+}}\,ds)\\ v(t)\leq C(\kappa+\frac{A_{2}}{\varepsilon_{2}^{+}}\int_{0}^{t}h(s)^{\varepsilon_{2}^{+}}v(s)^{1-\varepsilon_{2}^{+}}\,ds),\end{cases}$ (5.3)
@@
$x(t)=[(C^{+}\kappa)^{1/A_{1}}+C^{+}t]^{A_{1}},\quad y(t)=[(C^{+}\kappa)^{1/A_{2}}+C^{+}t]^{A_{2}},\qquad t\geq 0.$
@@
$\forall t\in[0,T],\quad\begin{cases}x(t)=C^{+}(\kappa+A_{1}\int_{0}^{t}(y(s))^{\varepsilon_{1}^{+}}\,ds)\\ y(t)=C^{+}(\kappa+A_{2}\int_{0}^{t}(x(s))^{\varepsilon_{2}^{+}}\,ds).\end{cases}$ (5.4)
@@
$h(t)<C(\kappa+A_{1}\int_{0}^{t}(y(s))^{\varepsilon_{1}^{+}}\,ds)<C^{+}(\kappa+A_{1}\int_{0}^{t}(y(s))^{\varepsilon_{1}^{+}}\,ds)=x(t)$
@@
$\forall t\in[0,T],\quad h(t)<[(C^{+}\kappa)^{1/A_{1}}+C^{+}t]^{A_{1}}\quad\hbox{and}\quad v(s)<[(C^{+}\kappa)^{1/A_{2}}+C^{+}t]^{A_{2}},$
@@
$x(t)=[(C^{+}\kappa)^{1/A_{1}}+C^{+}t)]^{A_{1}},\quad y(t)=[(C^{+}\kappa)^{1/B_{2}}+C^{+}t]^{B_{2}},\qquad t\geq 0,$
@@
$\forall t\in[0,T],\quad\begin{cases}x(t)=C^{+}(\kappa+A_{1}\int_{0}^{t}y(s)^{\varepsilon_{1}^{+}\varepsilon_{2}^{+}}\,ds)\\ y(t)=C^{+}(\kappa+B_{2}\int_{0}^{t}x(s)^{\varepsilon_{2}^{+}}y(s)^{1-\varepsilon_{2}^{+}}\,ds).\end{cases}$ (5.5)
@@ 5.4 Convergence of local times of a continuous-time simple random walk
for all $t,s\in[0,M]$ and $x,y\in[-M,M]$. Since $\tilde{\ell}^{T}_{t}(\cdot)$ is the piecewise linear interpolation of $\ell^{T}_{t}(\cdot)$, by a standard argument, equation (5.7) will follow upon proving
@@
Note that $\left(\frac{1}{\sqrt{T}}\xi(Tt)\right)_{t\geq 0}$ and $\left(\frac{1}{\sqrt{T}}M_{f_{\sqrt{T}x}}(Tt)\right)_{t\geq 0}$ are both martingales with quadratic variation bounded by $\frac{1}{T}Tt=t$, so, by the BDG inequality, both expectations above are bounded by $t^{p}$, up to a constant depending only on $p$. Therefore we obtain the following bound which is uniform in $x\in\frac{1}{\sqrt{T}}\mathbb{Z}$:
@@
where $(M_{f_{k,j}}(t))_{t\geq 0}$ is a martingale with quadratic variation $\langle M_{f_{k,j}},M_{f_{k,j}}\rangle_{t}\leq 2\int_{0}^{t}\mathbf{1}_{\{k\leq\xi(s)\leq j\}}\,ds$. We deduce that, for all $x,y\in\frac{1}{\sqrt{T}}\mathbb{Z}$ with $x<y$,
@@
where the martingale $M^{T}(t):=\frac{1}{\sqrt{T}}M_{f_{\sqrt{T}x,\sqrt{T}y}}(tT)$, $t\geq 0$, has quadratic variation
@@
Thus, using the triangle inequality once more, we get